# Do Construction Distributions Shape Formal Language Learning In German BabyLMs?

### v1

@@ Appendix B Excluded corpora
Appendix C Data cleaning
In line with best practices in language modeling, we extensively clean and normalize our data. Our cleaning script is available at [link removed for anonymization].
@@ Appendix C Data cleaning
Appendix D Exact construction proportions
@@
Appendix E Model hyperparameters and training details
@@
We reproduce the training and test loss curves for our models in Figure 3. For the test loss, we evaluated perplexity on a held-out, randomly sampled portion of each individual training corpus. We find no principal differences in loss development, although the character models and models trained on the cds data seem to converge the fastest. As the similar curves for train and test loss indicate, all models succeed in optimizing for their next-token prediction goal.
@@
Appendix F Repeated training runs
A common criticism in the BabyLM paradigm is the purported effect of training noise on model performance, which is hard to disentangle from real training data effects. While training and evaluating multiple random seeds for all our models would be too costly, we repeated two additional training runs for the character-level cds model with different random initializations (learning trajectories in Figure 4(a)) and two additional training runs where we re-sampled the cds dataset from our whole corpus with the exact same construction composition, but different content (learning curves in Figure 4(b)). In both cases, the learning trajectories do not differ tremendously. For the word-level phenomena (LexDec, Surprisal, AntiSurprisal), the curves overlap almost perfectly. For the syntax phenomena, we can see some variation and oscillation in the curves, but the trajectories still remain extremely similar (and do not differ in their steepness, the main effect that we see in Figure 2 between the datasets with different construction compositions).

### v2

@@ Appendix A Excluded corpora
Appendix B Data cleaning
In line with best practices in language modeling, we extensively clean and normalize our data.
@@ Appendix B Data cleaning
Appendix C Exact construction proportions
@@
Appendix D Model hyperparameters and training details
@@
We reproduce the training and test loss curves for our models in Figure 4. For the test loss, we evaluated perplexity on a held-out, randomly sampled portion of each individual training corpus. We find no principal differences in loss development, although the character models and models trained on the cds data seem to converge the fastest. As the similar curves for train and test loss indicate, all models succeed in optimizing for their next-token prediction goal. It should be noted that due to longer/shorter sequences in the different data mixtures and our choice of padding to the maximum sequence length, some models are trained for more steps, although the number of lexical tokens remains the same.
@@
Appendix E Repeated training runs
A common criticism towards the BabyLM paradigm is the purported effect of training noise on model performance, which is hard to disentangle from real training data effects. While training and evaluating multiple random seeds for all our models would be too costly, we repeated two additional training runs for the character-level cds model with different random initializations (learning trajectories in Figure 5(a)) and two additional training runs where we re-sampled the cds dataset from our whole corpus with the exact same construction composition, but different content (learning curves in Figure 5(b)). In both cases, the learning trajectories do not differ tremendously. For the word-level phenomena (LexDec, Surprisal, AntiSurprisal), the curves overlap almost perfectly. For the syntax phenomena, we can see some variation and oscillation in the curves, but the trajectories still remain extremely similar (and do not differ in their steepness, the main effect that we see in Figure 3 between the datasets with different construction compositions).